OpenStudy (anonymous):

can anyone help me with this? you don't have to answer for me, just help me through it please. the biomass of a deciduous forest is 50% carbon. additionally the biomass increases annually at a rate of 2.7x10^5 kg/hectare. calculate the mass of carbon accumulated and stored in 1.0 hectare if the forest in one year. answer to the nearest thousand kilograms

OpenStudy (whpalmer4):

The biomass increases by 2.7*10^5 kg/hectare each year. So 1.0 hectare of forest will accumulate 2.7*10^5 kg in one year. 50% of the biomass is carbon. Half of the accumulated biomass is therefore carbon. How much is 50% of 2.7*10^5 kg, rounded to the nearest 1000 kg?

OpenStudy (anonymous):

um if I did this right I got 135000 before rounding

OpenStudy (whpalmer4):

Yes, that is 50% of 2.7*10^5.

OpenStudy (anonymous):

is that all I had to do?

OpenStudy (whpalmer4):

Well, you don't know how much biomass was there originally, so you can't really do anything else! You can't figure out the total carbon stored in the hectare, just the amount added in one year.

OpenStudy (anonymous):

idk I typed it exactly as it said on my paper. I think they word it too complicated

OpenStudy (anonymous):

I mean I get mixed up on what to divide by what. that's my main problem

OpenStudy (whpalmer4):

you've got the figure for the biomass, and 50% of it is carbon. "x% of something" means multiply something by x% which is the same as multiplying by x/100. No division needed.

OpenStudy (anonymous):

well I just divided 2.7x10^5 by 2

OpenStudy (whpalmer4):

That is equivalent to taking 50%. But as you expressed confusion about "what to divide by what", I gave you a reliable procedure for calculating percentages. No extra charge :-)

OpenStudy (anonymous):

wow... I have a whole packet full of questions I don't know how to do :'(

OpenStudy (whpalmer4):

well, why don't you post a few of them and perhaps two sets of eyes will be better than one...

OpenStudy (anonymous):

I just don't want people to feel like I'm not trying

OpenStudy (anonymous):

do you want me to post them in the comments or close this question and open a new one

OpenStudy (whpalmer4):

Go ahead and post another one here...

OpenStudy (anonymous):

ok. well I tried this next one but it confused me too :/

OpenStudy (anonymous):

a population of glasswing butterfly exhibits logistic growth. the carring capacity of the population is 200 butterflies. and \[r_{\max}\] the maximum per capita growth rate of the population is 0.10 butterflies/ (butterflies x month). calculate the maximum population growth rate for the population if the maximum growth occurs when N=K/2. give answer to the nearest tenth

OpenStudy (whpalmer4):

Okay, doing a bit of reading...

OpenStudy (whpalmer4):

K is carrying capacity = 200 \(r_{max}\) is the (maximum) per capita growth rate = 0.10 N is population size = K/2

OpenStudy (anonymous):

take all the time you need. I'm also tweeting Zak Bagans so you don't have to reply right away. no hurry

OpenStudy (whpalmer4):

max population growth rate is \[\frac{dN}{dt} = rN(1-\frac{N}{K})= 0.10(K/2)(1-\frac{(K/2)}{K}) =\]

OpenStudy (whpalmer4):

Plug in the value of K (200) and I believe that is your answer. However, this would be the first such problem I've done, so my usual money-back guarantee is not in force :-)

OpenStudy (anonymous):

how in the heck did you get that

OpenStudy (anonymous):

I'm more confused now

OpenStudy (whpalmer4):

I went off and read the wikipedia page on logistic growth...

OpenStudy (anonymous):

we can't use wikipedia

OpenStudy (anonymous):

it might not be true

OpenStudy (whpalmer4):

I'm confident that you aren't expected to just know or invent these formulas, so perhaps doing the reading or viewing of your course materials again will reveal the source. Given that the section in question has bibliographic references to a number of texts and journal articles, I'm inclined to believe it until it is demonstrated I should not. Did your instructor provide a similar level of references for whatever he or she told you? :-)

OpenStudy (anonymous):

yeah there is no "course materials" it's literally a piece of paper with just these questions. that's all this packet is. no references or anything

OpenStudy (whpalmer4):

So, you just showed up, someone handed you a list of problems and said go do them, no lecturing, no reading, no nothing?

OpenStudy (whpalmer4):

You'll forgive my skepticism, but that's unlike any educational experience I've had in my many years...

OpenStudy (anonymous):

yes. it's for my AP bio class next year. summer assignment. I literally walked in and he said "here's your assignment, due the first day of school. have a good break"

OpenStudy (whpalmer4):

Perhaps you're expected to get the text to be used and read up about such things, or visit the library... Anyhow, looking at other sources on the internet, they seem to construct the same formula. https://www.youtube.com/watch?v=rXlyYFXyfIM http://personal.kenyon.edu/holdenerj/Math108Spring2007/classnotes/logisticgrowth/lesson4logisticgrowth.htm http://mathworld.wolfram.com/LogisticEquation.html https://home.comcast.net/~sharov/PopEcol/lec5/logist.html

OpenStudy (anonymous):

ok so what's the little d supposed to represent?

OpenStudy (whpalmer4):

oh, that's calculus notation, dN/dt is the rate of change of N with respect to t.

OpenStudy (whpalmer4):

my interpretation is that dN/dt is what they are asking for when they ask for the maximum population growth rate

OpenStudy (anonymous):

I have \[\frac{ dN }{ dt }=r _{\max}^N (\frac{ K-N }{ K })\]

OpenStudy (whpalmer4):

I could of course be incorrect. I'll ask my ex-girlfriend with the PhD in population biology if I run into her, but as we live on opposite sides of the continent, it's not very likely :-) Yes, that's the same. \[\frac{K-N}{K} = 1-\frac{N}{K}\]

OpenStudy (anonymous):

um I kinda need this done by next month

OpenStudy (anonymous):

how do I know what N is

OpenStudy (anonymous):

because K is the carrying capacity which is 200

OpenStudy (whpalmer4):

I think this is the key: "if the maximum growth occurs when N=K/2."

OpenStudy (anonymous):

but N is supposed to be the population size

OpenStudy (anonymous):

oh... hold on a sec

OpenStudy (whpalmer4):

Yes. Why can't it equal K/2? Btw, the youtube video I sent you looks pretty good.

OpenStudy (anonymous):

brain function in process... think it's registering

OpenStudy (anonymous):

so for this question the population is supposed to be 200/2 (K/2). so the maximum growth rate would be 10?

OpenStudy (whpalmer4):

I get 5, not 10.

OpenStudy (anonymous):

wait I did 100x .10. then went back and did \[\frac{ 200-100 }{ 100}\]x .10 and got .05

OpenStudy (whpalmer4):

\[\frac{dN}{dt} = rN(1-\frac{N}{K})= 0.10(K/2)(1-\frac{(K/2)}{K}) = 0.1(100)(1-\frac{1}{2}) = 5\]

OpenStudy (anonymous):

wtf is all that

OpenStudy (whpalmer4):

denominator of that fraction should be 200, not 100

OpenStudy (anonymous):

oops. my bad

OpenStudy (whpalmer4):

I think you're missing a factor of 100 in there somewhere, because I'm confident that the little butterflies are going to be able to add more than 5/100ths of a butterfly per month to the population

OpenStudy (anonymous):

I still got .05

OpenStudy (whpalmer4):

show me your formula and work. mine is right there, and correct, I believe.

OpenStudy (anonymous):

this is why I hate math

OpenStudy (anonymous):

I told you what I did

OpenStudy (whpalmer4):

no, you didn't. show me the entire formula you used, with the numbers you put in it, just like I did.

OpenStudy (anonymous):

I'm getting frustrated

OpenStudy (whpalmer4):

I haven't met anyone who doesn't at some point. How are you going to deal with it? Learning how to do so may be more valuable in the long run than any handful of facts you learn in AP biology.

OpenStudy (anonymous):

the question asks for the population growth and my paper gives me a different equation for that.

OpenStudy (whpalmer4):

Okay, what equation does it give?

OpenStudy (anonymous):

yeah I want to work delivering babies. I seriously doubt this stuff will help

OpenStudy (anonymous):

\[\frac{ dN }{ dt }= r_{\max}N \]

OpenStudy (anonymous):

but the dn/dt stuff doesn't matter cause it's set equal to rmaxN so as long as I have that it doesn't matter

OpenStudy (whpalmer4):

I don't think that equation shows logistic growth, from what I understand about it. Read the links I sent you, watch the video, make your own decision.

OpenStudy (anonymous):

but the question itself doesn't ask for logistic growth, it asks for population growth

OpenStudy (whpalmer4):

"a population of glasswing butterfly exhibits logistic growth"

OpenStudy (whpalmer4):

logistic growth is a particular kind of population growth

OpenStudy (whpalmer4):

In any case, it's time for me to go investigate the biology of eating a large cheeseburger :-) Good luck with the problems!

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

ooh a chili cheese burger sounds nice...

OpenStudy (anonymous):

I'm getting extremely pissed off right now.

OpenStudy (anonymous):

I got it